Grade 10Mechanics

A particle moving with simple harmonic motion has speeds V1 and V2 (V1>V2) and acceleration with magnitude F1 and F2 respectively at the points A and B which lie on the same side of the mean position o. sho w that AB=v1^2 -v2^2/ f1+f2

Mass murugan
9 Years agoGrade 10

1 Answer

Balaji
9 Years ago
Let OX =Y=A.sin(wt)differentiating, v= Awcos(wt)=w×sqrt(A^2-Y^2)Differentiating again , a= -Aw^2cos(wt)= -w^2 Y.Do you get the idea?